From GlueXWiki
Jump to: navigation, search

April 4, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
    • Package #3 repair
  2. Installation (Lubomir, Bill)
    • Cable installation: Rich's models [1], [2], [3]
    • Plans for practicing at Blue Crab
  3. Engineering (Bill)
  4. Electronics (Chris, Nick)
  5. Tests at EEL126 FDC E-log (Beni)
  6. Other


Participants: Fernando, Eugene, Bill, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave: will start wire deadening tomorrow; wire frames #2 and #3 will be the first to go with. In total 4 wire frames have to be deadened.

- Package #3 repair: we scanned all the wires of the HV sector #2 (on the wire frame of cell #6) using the camera of the motion station on the stringing table. There were three field wires with some stuff we could not clean. We disconnected these wires from the HV side, by un-soldering and cutting the end of the wire, but without success: the HV problem remained. Then we removed half of resistors at the HV side of the sense wires and the problem went away. Using this method we went down to 5 questionable wires, and at this moment Anatoly found a tip of a sense wire (~5mm) sticking out of the solder. After fixing this, the chamber now is holding HV already 10 hours. We will wait till tomorrow and then Casey will replace the three field wires, so the wire plane will be ready for installation by Monday. Bill suggested to glue/tape somehow the spacer ring to the wire frame when mounting it from the back of the package; will need more people for the cell isntallation.


- Yesterday, there was a "FDC cable installation" meeting on which Rich Getz presented his new plan shown on the three pictures attached above. Important change: the cables coming from the back of the installation cart don't have to run to the front of the magnet to reach the cable trays. This is possible because of the new openings in the floor next to the rails for the carts. These openings are just on the top of a new cable tray. To support the cables on the top, there will be a temporary movable frame. Beni suggested how to organize the cables on the top to make the insertion easier. Bill explained how the cables will be attached in between the packages and how the strain relief will be provided: using carbon tubes at the periphery.

- Everybody thinks this is a very good solution for the FDC cabling. Another important feature: the CDC cables will use a separate cable tray at the bottom below the platform. Thus, the installations of the FDC and CDC cables become independent.

- LV cabling: in the current scheme all the FDC LV cables (140) are connected to one rack. Fernando advocated to keep this scheme and bring half of the LV cables from one side to the other somehow at the front of the magnet. Doing this directly on the mesh might be problematic due to the small distance to BCAL, but Bill will look at this. Attaching additional support ring at the front of the magnet for the CDC cables (Bill) may help.

- We discussed possible practicing at Blue Crab with the following configuration. We will use two installation carts, going downstream, one for one half of the mesh cylinder, and the other for one FDC package and the mock-up. The idea is to prototype one quadrant of the signal cables (132). The cables for that quadrant will come from the FDC package and the rest will be attached to the mock-up representing the other three packages. The cables will be attached up to the end of the mesh running between the rails. We have all the major parts to do this. The bearings on the mesh have to be installed. Will move also 132 signal cables from 117 to Blue Crab.


- Bill made some modifications to the cooling system tests (see drawings attached), but the system will be ready very soon. He ordered also a sniffer for Fluorinert and as soon as we have it we can start the test. It will be done in 126, first just will check the pressure and make sure there are no leakages, then we will do the tests with the pre-amps powered.

- The tests with the pressurized Fluorinert in the cooling loops of package #4 will continue after we get the sniffer. Then we can continue mounting the manifolds on the other packages. Next week we will have the G10 pieces needed to install the manifolds at the 210deg point.


- Fernando: we have 20 fADC125 that can be used; still Cody has to finish with the tests. Generally there were problems with the quality of the boards, and the communication with the company wasn't easy. Now with the help of the procurement we hope they will improve the quality, but there will be a delay at least of a month.

Tests at EEL126

- Beni: there was an "accident" with the CAEN 1550P module: the HV on all the sense wire channels jumped from 2100V (nominal) to ~2350V (the hardware limit). Surprisingly, the chamber held this HV, except for a couple of channels. Thus (unwillingly) we proved that the plateau of the chamber can be as wide as 350V, and tested all the HV capacitors at once! Beni replaced the HV module with a new one and all the channels were OK. This is the second CAEN module showing this problem. Therefore, one has to always lower the hardware limit (there's a pot on the back) and don't rely on the software only.

- Beni successfully tested his code running on the ROCs that calculates the charge and the time using the raw data from the fADC125s. However, this introduces some noticeable dead time in the DAQ. This might be important when reading 14 fADCs in the full electronics test, instead of one per crate as it is now. Simon suggested looking at the code and try to speed it up. Beni also may give up the precise time calculations (using Gerard's algorithm) that we don't need now.

- Beni is done with package #2 and we can move it back to Blue Crab.